# A mechanism for ice growth on the surface of a spherical water droplet

### v1

@@ III Casimir-Lifshitz free energy for the concentric sphere geometry
in which $\kappa_{i;n}=2\pi nT\sqrt{\varepsilon_{i}\mu_{i}},\ i=1,2,3$ (with the indices corresponding to water, ice, vapor respectively), the temperature $T=273.16\rm\ K$, and the bracket $[\cdot,\cdot]_{\varepsilon}$, involving the permittivity $\varepsilon$, is defined as
$[f(x),g(x)]_{\varepsilon}=\frac{f^{\prime}(x)}{\varepsilon_{f}}g(x)-f(x)\frac{g^{\prime}(x)}{\varepsilon_{g}}.$ (3)
@@
Suppose the separation between inner and outer interfaces $d=b-a$ goes to zero with $a$ remaining constant. Then in this limit, the retardation effects are too trivial to be included, and only non-retarded contributions survive. The TM CL energy, as an instance, in this case is
@@
Even if the retardation is fully counted, the geometry of the configuration acts as another factor making the situation more complicated. In contrast with the small-separation limit above, if $d=b-a$ kept constant but $a\rightarrow\infty$ (referred to as the “planar limit” here), then TM Casimir-Lifshitz free energy $F_{\rm H}$, based on the Appendix, behaves as
@@
×$\times\frac{\varepsilon_{1}\widetilde{\kappa}_{2,n}-\varepsilon_{2}\widetilde{\kappa}_{1,n}}{\varepsilon_{1}\widetilde{\kappa}_{2,n}+\varepsilon_{2}\widetilde{\kappa}_{1,n}}e^{-2\widetilde{\kappa}_{2,n}d}\bigg{)},$
@@ IV Results and Discussions
[…]
@@
[…]
10 1415 $-1.38\times 10^{-5}$ $2.89\times 10^{8}$
20 1405 $-2.75\times 10^{-5}$ $3.62\times 10^{7}$
40 1386 $-5.50\times 10^{-5}$ $4.53\times 10^{6}$
60 1368 $-8.23\times 10^{-5}$ $1.35\times 10^{6}$
80 1351 $-1.09\times 10^{-4}$ $5.72\times 10^{5}$
100 1334 $-1.36\times 10^{-4}$ $2.94\times 10^{5}$
200 1262 $-2.66\times 10^{-4}$ $3.90\times 10^{4}$
400 1163 $-4.88\times 10^{-4}$ $5.87\times 10^{3}$
600 1106 $-6.51\times 10^{-4}$ $2.20\times 10^{3}$
800 1072 $-7.61\times 10^{-4}$ $1.18\times 10^{3}$
1700 1025.18 $-9.31\times 10^{-4}$ $3.12\times 10^{2}$
2500 1020.18 $-9.34\times 10^{-4}$ $1.79\times 10^{2}$
3000 1020.39 $-9.20\times 10^{-4}$ $1.41\times 10^{2}$
3500 1021.32 $-9.04\times 10^{-4}$ $1.16\times 10^{2}$
4000 1022.52 $-8.88\times 10^{-4}$ $9.80\times 10^{1}$
5000 1025.26 $-8.58\times 10^{-4}$ $7.50\times 10^{1}$
5500 1026.14 $-8.45\times 10^{-4}$ $6.71\times 10^{1}$
$\infty$ 1043.84 $-6.19\times 10^{-4}$
@@ V Conclusions
Suppose the separation between the inner and outer interface $d=b-a$ goes to zero, while $a$ is kept constant. Obviously, in the zero-separation case, $d=0$, the Casimir free energy in the concentric configuration diverges, as the corresponding case in the planar system. To capture its small-separation behaviors, we consider the separation nonzero, but infinitely small, that is, the TM Casimir free energy per unit area with respect to the inner interface, for instance (TE part following the same arguments), expressed as
@@
which result in the $-r_{32}^{\rm H}(n,l;b)r_{21}^{\rm H}(n,l;a)$ with $d\rightarrow 0_{+}$ expressed as
@@
On the other way round, when $d=b-a$ kept constant but $a\rightarrow\infty$, we arrive at another limitting case (or referred to as the “planar limit”). Surprisingly, the UAE is also appropriate to be employed to investigate behaviors in this limit Parashar et al. (2019). Then the arguments above can be directly followed, except for the term $-r_{32}^{\rm H}(n,l;b)r_{21}^{\rm H}(n,l;a)$ in the instance above, which can be calculated, given $a\rightarrow\infty$, as
@@
×$\quad\times\frac{\varepsilon_{1}(1+z_{2,a}^{2})^{1/2}-\varepsilon_{2}(1+z_{1,a}^{2})^{1/2}}{\varepsilon_{1}(1+z_{2,a}^{2})^{1/2}+\varepsilon_{1}(1+z_{1,a}^{2})^{1/2}}$
@@
×$\times\frac{\varepsilon_{1}\sqrt{\nu^{2}/a^{2}+\varepsilon_{2}\zeta_{n}^{2}}-\varepsilon_{2}\sqrt{\nu^{2}/a^{2}+\varepsilon_{1}\zeta_{n}^{2}}}{\varepsilon_{1}\sqrt{\nu^{2}/a^{2}+\varepsilon_{2}\zeta_{n}^{2}}+\varepsilon_{1}\sqrt{\nu^{2}/a^{2}+\varepsilon_{1}\zeta_{n}^{2}}}.$
@@
×$\times\frac{\varepsilon_{1}\widetilde{\kappa}_{2,n}-\varepsilon_{2}\widetilde{\kappa}_{1,n}}{\varepsilon_{1}\widetilde{\kappa}_{2,n}+\varepsilon_{2}\widetilde{\kappa}_{1,n}}e^{-2\widetilde{\kappa}_{2,n}d}\bigg{)},$
@@
with $+$ and -- signifying the solution being 00 and finite at the infinity and origin, respectively. Within a uniform background, the solutions are denoted as $\mathfrak{f}_{\pm}(\zeta,l;r)=\mathfrak{e}_{\pm}(\zeta,l;\kappa r)=\mathfrak{h}_{\pm}(\zeta,l;\kappa r)$, and $\kappa=\sqrt{\varepsilon\mu}|\zeta|$. Then, for a layered concentric system with the permittivity and permeability being
$(\varepsilon,\mu)(r)=\left\{\begin{array}[]{cc}(\varepsilon_{1},\mu_{1}),&r<a_{1},\\ (\varepsilon_{2},\mu_{2}),&a_{1}<r<a_{2},\\ \vdots&\vdots\\ (\varepsilon_{N},\mu_{N}),&a_{N-1}<r<a_{N},\\ (\varepsilon_{N+1},\mu_{N+1}),&r>a_{N},\\ \end{array}\right.$ (29)
@@
$\mathfrak{e}_{+}(r)=\left\{\begin{array}[]{cc}A_{1}\mathfrak{f}_{1,+}(r)+B_{1}\mathfrak{f}_{1,-}(r),&r<a_{1},\\ A_{2}\mathfrak{f}_{2,+}(r)+B_{2}\mathfrak{f}_{2,-}(r),&a_{1}<r<a_{2},\\ \vdots&\vdots\\ A_{N}\mathfrak{f}_{N,+}(r)+B_{N}\mathfrak{f}_{N,-}(r),&a_{N-1}<r<a_{N},\\ \mathfrak{f}_{N+1,+}(r),&r>a_{N},\\ \end{array}\right.\mathfrak{e}_{-}(r)=\left\{\begin{array}[]{cc}\mathfrak{f}_{1,-}(r),&r<a_{1},\\ C_{2}\mathfrak{f}_{2,+}(r)+D_{2}\mathfrak{f}_{2,-}(r),&a_{1}<r<a_{2},\\ \vdots&\vdots\\ C_{N}\mathfrak{f}_{N,+}(r)+D_{N}\mathfrak{f}_{N,-}(r),&a_{N-1}<r<a_{N},\\ C_{N+1}\mathfrak{f}_{N+1,+}(r)+D_{N+1}\mathfrak{f}_{N+1,-}(r),&r>a_{N},\\ \end{array}\right.$ (40)
@@ Acknowledgements.
We thank the Terahertz Physics and Devices Group, Nanchang University for the strong computational facility support. Y.L. was supported by the National Natural Science Foundation of China (grant number 12304396). P.P. thanks Thushari Jayasekara for collaborative assistance during the course of this work. The research by O.I.M. and M.B. is part of the project No. 2022/47/P/ST3/01236 co-funded by the National Science Centre and the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 945339. The research by M.B. took place at the "ENSEMBLE3 - Centre of Excellence for nanophotonics, advanced materials and novel crystal growth-based technologies" project (GA No. MAB/2020/14) carried out within the International Research Agendas programme of the Foundation for Polish Science co-financed by the European Union under the European Regional Development Fund, the European Union’s Horizon 2020 research and innovation programme Teaming for Excellence (GA. No. 857543). MB’s research contributions to this publication were created as part of the project of the Minister of Science and Higher Education "Support for the activities of Centers of Excellence established in Poland under the Horizon 2020 program" under contract No. MEiN/2023/DIR/3797. We gratefully acknowledge Poland’s high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016228. C.P. thanks the European Union’s Horizon 2020 research and innovation programme (GA. No. 101058694).

### v2

# arXiv:2502.13155v2[cond-mat.mtrl-sci] 28 Feb 2026

@@ III Casimir-Lifshitz free energy for the concentric sphere geometry
in which $\kappa_{i;n}=2\pi nT\sqrt{\varepsilon_{i}\mu_{i}},\ i=1,2,3$ (with the indices corresponding to water, ice, vapor respectively), the temperature $T=273.16\rm\ K$, and the bracket $\cdot,\cdot]_{\varepsilon}$, involving the permittivity $\varepsilon$, is defined as
$f(x),g(x)]_{\varepsilon}=\frac{f^{\prime}(x)}{\varepsilon_{f}}g(x)-f(x)\frac{g^{\prime}(x)}{\varepsilon_{g}}.$ (3)
@@
Suppose the separation between inner and outer interfaces d=bad=b-a goes to zero with $a$ remaining constant. Then in this limit, the retardation effects are too trivial to be included, and only non-retarded contributions survive. The TM CL energy, as an instance, in this case is
@@
Even if the retardation is fully counted, the geometry of the configuration acts as another factor making the situation more complicated. In contrast with the small-separation limit above, if d=bad=b-a kept constant but $a\rightarrow\infty$ (referred to as the “planar limit” here), then TM Casimir-Lifshitz free energy $F_{\rm H}$, based on the Appendix, behaves as
@@
$\times\frac{\varepsilon_{1}\widetilde{\kappa}_{2,n}-\varepsilon_{2}\widetilde{\kappa}_{1,n}}{\varepsilon_{1}\widetilde{\kappa}_{2,n}+\varepsilon_{2}\widetilde{\kappa}_{1,n}}e^{-2\widetilde{\kappa}_{2,n}d}\bigg),$
@@ IV Results and Discussions
It was recently predicted that minimization of Casimir-Lifshitz free energies could induce ice layer formation at the surfaces of water Fiedler et al. (2020); Luengo-Márquez and MacDowell (2021); Luengo-Marquez et al. (2022), as well as on AgI particles in clouds Luengo-Márquez and MacDowell (2021). The mechanism for ice formation has also been invoked for self-preservation of gas hydrates Boström et al. (2021). However, all those works, predicting the minimization of free energy with the ice growth, exploited theories with the planar approximation, which in the submicron size, are often better described as concentric spherical systems. The study on energies in the actual submicron concentric spherical systems unveils interesting new physics. The CL free energy curves as functions of the ice layer thickness, for water droplets with given inner radii (a=10, 50a=10,\ 50, and $100\rm\ nm$), are shown in Fig. 1b. Two things are of immediate interest. Firstly, the depth of the minimum for the free energy per unit area increases, as the water droplet radius increases. It follows from energy considerations that close to micron-thick ice layers grow on nano-sized water droplets under the right conditions. Thereafter, the inner liquid core of the layered droplet grows at the expense of the inner parts of the ice coating. The ice film thickness, in turn, is kept close to constant by water vapor adsorption, leading to an expanding inner radius ($a$) and thus the radius of the composite nuclei ($b$). Notably, with a thinner ice layer, it is not energetically favorable for the water droplet to grow since the energy increase with increasing $a$ for small ice layer thicknesses. The second important observation is that the relative amount of ice in the droplet, which minimizes the energy, increases with decreasing water droplet radius. As can be seen in Table 1, the relative amount of ice growth is predicted to be dramatic for nano-sized water droplets. This leads us ultimately to predict an enhanced ice growth mechanism for ice nucleating nano-droplets. This occurs in a new “secondary” way, via a reduced CL free energy, that leads to the simultaneous growth of both the interior water droplet and the ice coating with the adsorption of water molecules from the vapor phase. This results in larger volumes of ice, but should not be mixed up with the traditional secondary ice processes. Furthermore, the planar approximation, compared with a water droplet of a radius 10 nm, gives a correct order of magnitude estimate for the predicted ice coating thickness. It should be emphasized that for a fixed water droplet size, the repulsive Casimir-Lifshitz stress can be substantially increased below the equilibrium ice film thickness. This can lead to the establishment of the initial growth of an ice layer due to dispersion energies.
@@
$a$ (nm) ice thickness (nm) $F_{\rm min}$ (nJ/m2) volume ratio ($\%$)
10 1415 $1.38\times 10^{-5}$ $2.89\times 10^{8}$
20 1405 $2.75\times 10^{-5}$ $3.62\times 10^{7}$
40 1386 $5.50\times 10^{-5}$ $4.53\times 10^{6}$
60 1368 $8.23\times 10^{-5}$ $1.35\times 10^{6}$
80 1351 $1.09\times 10^{-4}$ $5.72\times 10^{5}$
100 1334 $1.36\times 10^{-4}$ $2.94\times 10^{5}$
200 1262 $2.66\times 10^{-4}$ $3.90\times 10^{4}$
400 1163 $4.88\times 10^{-4}$ $5.87\times 10^{3}$
600 1106 $6.51\times 10^{-4}$ $2.20\times 10^{3}$
800 1072 $7.61\times 10^{-4}$ $1.18\times 10^{3}$
1700 1025.18 $9.31\times 10^{-4}$ $3.12\times 10^{2}$
2500 1020.18 $9.34\times 10^{-4}$ $1.79\times 10^{2}$
3000 1020.39 $9.20\times 10^{-4}$ $1.41\times 10^{2}$
3500 1021.32 $9.04\times 10^{-4}$ $1.16\times 10^{2}$
4000 1022.52 $8.88\times 10^{-4}$ $9.80\times 10^{1}$
5000 1025.26 $8.58\times 10^{-4}$ $7.50\times 10^{1}$
5500 1026.14 $8.45\times 10^{-4}$ $6.71\times 10^{1}$
$\infty$ 1043.84 $6.19\times 10^{-4}$
@@ V Conclusions
Suppose the separation between the inner and outer interface d=bad=b-a goes to zero, while $a$ is kept constant. Obviously, in the zero-separation case, $d=0$, the Casimir free energy in the concentric configuration diverges, as the corresponding case in the planar system. To capture its small-separation behaviors, we consider the separation nonzero, but infinitely small, that is, the TM Casimir free energy per unit area with respect to the inner interface, for instance (TE part following the same arguments), expressed as
@@
which result in the $r_{32}^{\rm H}(n,l;b)r_{21}^{\rm H}(n,l;a)$ with $d\rightarrow 0_{+}$ expressed as
@@
On the other way round, when d=bad=b-a kept constant but $a\rightarrow\infty$, we arrive at another limitting case (or referred to as the “planar limit”). Surprisingly, the UAE is also appropriate to be employed to investigate behaviors in this limit Parashar et al. (2019). Then the arguments above can be directly followed, except for the term $r_{32}^{\rm H}(n,l;b)r_{21}^{\rm H}(n,l;a)$ in the instance above, which can be calculated, given $a\rightarrow\infty$, as
@@
$\quad\times\frac{\varepsilon_{1}(1+z_{2,a}^{2})^{1/2}-\varepsilon_{2}(1+z_{1,a}^{2})^{1/2}}{\varepsilon_{1}(1+z_{2,a}^{2})^{1/2}+\varepsilon_{1}(1+z_{1,a}^{2})^{1/2}}$
@@
$\times\frac{\varepsilon_{1}\sqrt{\nu^{2}/a^{2}+\varepsilon_{2}\zeta_{n}^{2}}-\varepsilon_{2}\sqrt{\nu^{2}/a^{2}+\varepsilon_{1}\zeta_{n}^{2}}}{\varepsilon_{1}\sqrt{\nu^{2}/a^{2}+\varepsilon_{2}\zeta_{n}^{2}}+\varepsilon_{1}\sqrt{\nu^{2}/a^{2}+\varepsilon_{1}\zeta_{n}^{2}}}.$
@@
$\times\frac{\varepsilon_{1}\widetilde{\kappa}_{2,n}-\varepsilon_{2}\widetilde{\kappa}_{1,n}}{\varepsilon_{1}\widetilde{\kappa}_{2,n}+\varepsilon_{2}\widetilde{\kappa}_{1,n}}e^{-2\widetilde{\kappa}_{2,n}d}\bigg),$
@@
with ++ and - signifying the solution being 0 and finite at the infinity and origin, respectively. Within a uniform background, the solutions are denoted as $\mathfrak{f}_{\pm}(\zeta,l;r)=\mathfrak{e}_{\pm}(\zeta,l;\kappa r)=\mathfrak{h}_{\pm}(\zeta,l;\kappa r)$, and $\kappa=\sqrt{\varepsilon\mu}|\zeta|$. Then, for a layered concentric system with the permittivity and permeability being
[…]
@@
[…]
@@ Acknowledgements.
We thank the Terahertz Physics and Devices Group, Nanchang University for the strong computational facility support. Y.L. was supported by the National Natural Science Foundation of China (Grant No. 12304396). P.P. thanks T. Jayasekara for collaborative assistance during the course of this work. The research by O.I.M. and M.B. is part of Project No. 2022/47/P/ST3/01236 cofunded by the National Science Centre and the European Union’s Horizon 2020 research and innovation program under the Marie Sk?odowska-Curie Grant Agreement No. 945339. Institutional and infrastructural support for the ENSEMBLE3 Centre of Excellence was provided through the ENSEMBLE3 project (MAB/2020/14) delivered within the Foundation for Polish Science International Research Agenda Programme and co-financed by the European Regional Development Fund and the Horizon 2020 Teaming for Excellence initiative (Grant Agreement No. 857543), as well as the Ministry of Education and Science initiative “Support for Centres of Excellence in Poland under Horizon 2020”. We gratefully acknowledge Poland’s high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) for providing computer facilities and support within computational Grant No. PLG/2023/016228. C.P. thanks the European Union¡¯s Horizon 2020 research and innovation program (GA No. 101058694).